im brandon, the question i have is for algebra A. Its solving systems using substitution, problem is Y=x-4 and 2x-5y=2 . can you explain how to do this .
Answer by mananth(16946)

y = x-4
2 x + -5 y = 2
Substitute y in the second equation
2 x + -5 (x-4) = 2
2 x -5x+20 = 2
2 x -5 x = -18
-3 x = -18
/ -3
x= 6
Plug the value of x in (1)
-1 x + 1 y = -4
-1 * 6 + 1 y = -4
-6 + 1 y = -4
y=6-4
y = 2

m.ananth@hotmail.ca